OpenStudy (djbreezy):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x + @MinecraftNoob

OpenStudy (djbreezy):

@Sam_Aka_Sara u know how to do this??

OpenStudy (djbreezy):

@Nnesha

OpenStudy (djbreezy):

@Black_ninja123

OpenStudy (djbreezy):

@Nnesha

OpenStudy (djbreezy):

i need help like Nowwwwww

OpenStudy (anonymous):

oh sorry idk how to do this i need ur help!

OpenStudy (djbreezy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you will use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

line 1 has the two points (0,3) and (1,5) line 2 has the two points (-1,1) and (-2,-1)

OpenStudy (djbreezy):

i need the slope

OpenStudy (djbreezy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

does that formula look familiar?

OpenStudy (djbreezy):

i only can put 2 numbers no comma

OpenStudy (djbreezy):

@peachpi

jimthompson5910 (jim_thompson5910):

you subtract the y values then subtract the x values finally, divide the differences (y over x)

jimthompson5910 (jim_thompson5910):

for line 1 ydifference: 5-3 = 2 xdifference: 1-0 = 1 divide the differences (y/x) to get 2/1 = 2 so the slope of line 1 is m = 2. Does that make sense?